x^2-8x+9=0 solve by "completing the square"
Answer by richard1234(7193)

We want to write x^2 - 8x + 9 as (x-4)(x-4) + C, where C is a constant. The constant term of (x-4)(x-4) is 16, so if we add 7 from both sides of our equation we get

x^2 - 8x + 16 = 7

(x-4)^2 = 7

x-4 = +/- sqrt(7)

x = 4 +/- sqrt(7)